OpenStudy (anonymous):

Please help with this problem? the range of a projectile fired at an angle θ with the horizontal and with the initial velocity of v0 feet per second is r=1/32v0^2 sin2θ where r is measured in feet. A golfer strikes a golfball at 115 feet per second. Ignoring the effects of air resistance, at what angle must the golfer hit the ball so that it travels 160 feet?

OpenStudy (anonymous):

11.55 degrees

OpenStudy (anonymous):

if taken 9.8 as g else if ten then it will be 12

OpenStudy (anonymous):

@harpreetsk ANSWER CHOICES: 23 degrees 11 degrees 17 degrees 7 degrees 26 degrees

OpenStudy (anonymous):

11 is correct then it would be result if taken g as 9.62

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

the way to do this is put in the numbers they give you into the equation \[ r=\frac{v_0^2}{32} \sin2θ \] r , the distance is 160 ft the velocity V0 is 115 \[ 160=\frac{115^2}{32} \sin2θ \] multiply both sides by the inverse (flipped fraction) \[ 160 \cdot \frac{32}{115^2} = \sin 2\theta \] you will need a calculator to find the inverse sin of both sides that will give you 2θ divide by 2 to get θ